OpenStudy (anonymous):

The function f(x)=x^2. the graph of g(x) is f(x) translated to the right 3 units and down 3 units. What is the function rule for g(x)?

OpenStudy (whpalmer4):

to move a function up or down is easy — you just add the appropriate value to the result of the function. to move down 3 units, we add -3 to f(x).

OpenStudy (whpalmer4):

to move right or left, we have to adjust the number we put into the function.

OpenStudy (anonymous):

Up and down you add to the equation Left and right you you add to the exponent.

OpenStudy (whpalmer4):

@nelsonjedi not the exponent, the function argument

OpenStudy (anonymous):

So is the function rule the equation? or is it the rules of moving the equation?

OpenStudy (whpalmer4):

so if we add 3 to the argument of the function, the result will be the same as if we had computed the function 3 units to the right. for example, f(0+3) = f(3), so whatever value the function normally has at x = 3 will now appear at x = 0. does that make sense?

OpenStudy (anonymous):

Wouldn't it be the other way around? whatever appears at x=0 would appear at x=3

OpenStudy (anonymous):

Sorry my bad. Yes the function arguement

OpenStudy (whpalmer4):

adding to the function argument shifts the graph to the left. subtracting from the function argument shifts the graph to the right.

OpenStudy (whpalmer4):

No, make yourself a graph and try it. use f(x) = x^2. here's a graph of x^2 vs (x+3)^2

OpenStudy (whpalmer4):

the blue graph is x^2, the purple is (x+3)^2. at x = -3, the purple graph = 0, which is what the blue graph = at x = 0.

OpenStudy (anonymous):

Okay. so would my answer be g(x)=3x^2-3

OpenStudy (whpalmer4):

so if we wanted that same parabola to have its vertex over at (5,0) instead of the (0,0) that we get with \(f(x) = x^2\) we have to subtract 5 from the argument to \(f(x)\), right?

OpenStudy (anonymous):

or g(x)=3+x^2-3

OpenStudy (whpalmer4):

No. You've got the up/down part correct, but multiplying x^2 by 3 just changes the scale up and down... This graph shows the original in blue, and the shift down three units (gotten by adding -3 to the result of f(x))

OpenStudy (whpalmer4):

Like I said, that part is easy. If we want g(x) to just be a shift f(x) down 3 units, we just say \[g(x) = f(x)-3\]

OpenStudy (whpalmer4):

To shift f(x) right or left, we have to do something like \[g(x) = f(x-a)\]

OpenStudy (whpalmer4):

and to do both right/left and up/down, we'll have something like \[g(x) = f(x-a) + b\]

OpenStudy (whpalmer4):

because \(f(x) = x^2\] that will become \[g(x) = (x-a)^2+b\]

OpenStudy (anonymous):

so the equation would be g(x)=f(x-3)+3 or g(x)=f(x-3)-3

OpenStudy (whpalmer4):

the latter one, yes.

OpenStudy (anonymous):

Okay. thanks

OpenStudy (whpalmer4):

sorry about messing up the formatting a post ago, that should have been because \(f(x) = x^2\) that will become \[g(x) = (x-a)^2+b\] (a and b can be negative or positive, of course)

OpenStudy (anonymous):

And that would be my answer to the original question?

OpenStudy (whpalmer4):

well, probably you should expand it out. \[g(x) = f(x-3)-3 = \]can you expand that, given that \(f(x) = x^2\)

OpenStudy (anonymous):

expand it how?

OpenStudy (whpalmer4):

well, replace f(x-3) with its expanded form: how would you compute the value of f(x-3)?

OpenStudy (whpalmer4):

what you have now is in fact a correct answer, but you should understand how to complete the process because that may be what is desired here

OpenStudy (whpalmer4):

what you've written is math-ese for "make g(x) by translating the graph of f(x) to the right 3 units and down 3 units" but it doesn't tell us what g(x) actually is so that we could graph it, do you see what I mean?

OpenStudy (anonymous):

distribute the f into the parenthesis?

OpenStudy (whpalmer4):

if \[f(x) = x^2\]\[f(x+1) = (x+1)^2\]right?

OpenStudy (whpalmer4):

just like \[f(3) = (3)^2\]

OpenStudy (anonymous):

Yes

OpenStudy (whpalmer4):

okay, so if \(g(x) = f(x-3)-3\), what does \(g(x)=\) in terms of \(x\), no \(f()\) allowed?

OpenStudy (anonymous):

It would be the some problem without the f, right?

OpenStudy (whpalmer4):

\[g(x) = f(x-3)-3\]replace \( f(x-3)\) with \((x-3)^2\) \[g(x) = f(x-3)-3 = (x-3)^2-3\]and now expand that out

OpenStudy (anonymous):

Expand it?

OpenStudy (whpalmer4):

multiply it out

OpenStudy (whpalmer4):

\[g(x) = (x-3)^2-3 =(x-3)(x-3)-3 = x^2-3x-3x+9-3\]\[ = x^2-6x+6\]

OpenStudy (whpalmer4):

now we have an equation for \(g(x)\) that we can use without needing to know anything about \(f(x)\)

OpenStudy (whpalmer4):

but if you plug in a bunch of values and plot it, you'll see that in fact, it is the graph of f(x) shifted 3 units to the right and 3 units down

OpenStudy (anonymous):

Wow, okay. I didn't know you could do that just by multiplying it out. Thank you.

OpenStudy (whpalmer4):

Yes, \[(x-3)^2 = (x-3)(x-3)\]

OpenStudy (anonymous):

Right.

OpenStudy (whpalmer4):

think of the \(f(x)\) notation as a recipe for how to construct the function. if you have something like \(f(x) = x^3+2x^2+5x-3\) that means you can construct the formula for \(f(a)\) by cubing a, adding 2a^2, adding 5a, and subtracting 3, for whatever value a is. if a happens to be some ugly polynomial with 17 terms, well, it's going to be a lot of work, but the approach is the same :-)

OpenStudy (anonymous):

Okay. I wish I could look at math like that. I can never seem to see how that pieces fit together. Thanks. :D

OpenStudy (whpalmer4):

and the \[g(x) = f(x-3) -3\]notation works for any function \(f(x)\)! Say we had a different parabola, \(f(x) = x^2+2x-4\) If we wanted to shift it 3 left and 3 down, \[g(x) = f(x-3) -3 = (x-3)^2+2(x-3)-4 -3\]\[=x^2-6x+9+2x-6-4-3 = x^2-4x-4\]\[g(x) = x^2-4x-4\] And here they both are, graphed together:

OpenStudy (whpalmer4):

anyhow, the key takeaway from this exercise is: to shift a function up or down by \(a\), add or subtract from the function result: \[f(x) \pm a\] to shift a function left or right by \(a\), add or subtract from the function argument: \[f(x\pm a)\]

OpenStudy (anonymous):

Okay. I got it, I think. haha. thanks

OpenStudy (whpalmer4):

you'll have many opportunities to test your understanding of this in the coming years :-)

OpenStudy (anonymous):

I hope not too much in college.